OpenStudy (blondey_315):

A scale drawing has a scale of 1.75 inches=1.5 feet. How long is a wall that measures 14 inches in the scale drawing?

OpenStudy (anonymous):

12 feet

OpenStudy (anonymous):

12

OpenStudy (blondey_315):

dankashen

OpenStudy (mathstudent55):

Use a proportion: \(\dfrac{1.75~in.}{1.5~ft} = \dfrac{14~in.}{x} \) Now cross multiply to find the answer: \(1.75x = 14 \times 1.5\) 1.75x = 21 x = 12